OpenStudy (ujjwal):

\[f(x)=\left\{\begin{matrix} sinx:x\neq n\pi ,n\in \mathbb{Z}\\0,otherwise & \end{matrix}\right.\ g(x)=\left\{\begin{matrix} x^2+1, if x\neq 0,2\\x, if x=0\\5, if x=2 \end{matrix}\right.\] then \[\lim_{x \to 0}g[f(x)]\]=

OpenStudy (experimentx):

Hmm ... I think you can do this \[ \lim_{x \rightarrow 0} f(g(x)) = \lim_{x \rightarrow 0} f\left ( \lim_{x \rightarrow 0}g(x) \right )\]

OpenStudy (ujjwal):

And how do we know if \(x^2\)+1 is equal to or not equal to n\(\pi\) ?

OpenStudy (ujjwal):

OS... tyestai ho.. F5 helps.

OpenStudy (experimentx):

Hmm ... g(x) = x^2 + 1 as x->0 which tends to 1

OpenStudy (experimentx):

which is \( \neq \) nπ

OpenStudy (experimentx):

still ... i'm not so sure ...

OpenStudy (anonymous):

Isnt it g(f(x))?

OpenStudy (ujjwal):

Ah! lets begin again.

OpenStudy (experimentx):

Oh ... my bad ;(

OpenStudy (experimentx):

i guess it would be zero in both cases

OpenStudy (experimentx):

for f(x)

OpenStudy (anonymous):

My bad x NOT equal to npi. :/

OpenStudy (anonymous):

So its, \[\lim_{x \rightarrow 0} \sin^2 x + 1\] Which is basically 1.

OpenStudy (ujjwal):

Here x tends to 0 implies x tends to 0\(\pi\) and is not equal to n\(\pi\) so, f(x)= sinx and then g(sinx)= \(\sin^2 x\)+1. =1 (putting x=0) Is that correct?

OpenStudy (anonymous):

Yesss. :)

OpenStudy (ujjwal):

thanks!